OpenStudy (anonymous):

What is the result of isolating x2 in the equation below? x2 + (y - 5)2 = 30 A. x2 = -y2 + 10y + 25 B. x2 = -y2 + 10y + 5 C. x2 = y2 + 10y + 5 D. x2 = 30 - y2

OpenStudy (anonymous):

Is the (y-5) squared or *2?

OpenStudy (anonymous):

squared

OpenStudy (anonymous):

Foil (y-5)(y-5)= y^2 -5y-5y+25 did you get how to do that?

OpenStudy (anonymous):

would you mind just telling me the answer i need to get going soon

OpenStudy (anonymous):

x^2-10y+25=30 x^2-10y=5 x^2=10y+5

OpenStudy (anonymous):

oops i miss the -y^2 it x^2=-y^2+10y+5